Dephart produces exholters. Variable is applied based on direct labor hours.

Dephart expects production for the coming period to be 4560 units. At that level of activity, variable overhead is expected to be $9120. It should take 2 direct labor hours to produce 1 exholter.

Actual results are as follows:
Number of exholters produced: 5700
Actual direct labor hours: 13680
Actual variable overhead cost: $10944


1. Complete the following standard cost sheet:

Quantity

@

Price

Total

Direct materials

7 pieces

$3.00

$21.00

Direct labor

2 hours

$10.00

$20.00

Variable Overhead

2 hours

$

$

Standard Unit Cost

$



What would be in a flexible budget for materials cost?

What would be in a flexible budget for labor cost?

What would be in a flexible budget for variable overhead cost?  
Answer #1

Solution 1:

Budgeted Direct labor hours = 4560*2 = 9120 hours

Variable overhead rate per hour = $9120 / 9120 = $1 per hour

Quantity @ Price Total
Direct materials 7 Pieces $3.00 $21.00
Direct labor 2 hours $10.00 $20.00
Variable Overhead 2 hours $1.00 $2.00
Standard Unit Cost $43.00

Flexible budget for materials cost = Actual unit * Standard Material cost per unit = 5700 * $21 = $119,700

Flexible budget for labor cost = Actual unit * Standard Labor cost per unit = 5700 * $20 = $114,000

Flexible budget for Variable overhead cost = Actual unit * Standard variable Overhead cost per unit = 5700 * $2 = $11,400